# Data on the Brain & Mind

### v1

@@ 1 Introduction
A central task of the brain is to process sensory information in order to guide adaptive behavior, often studied through perceptual decision-making. This process is shaped by several biological constraints. Neurons communicate through discrete action potentials (or spikes), the nervous system is subject to noise at multiple levels [faisal2008noise], and operates under limited metabolic resources [laughlin1998metabolic, laughlin2001energy, stone2018principles]. Together, these factors impose fundamental limits on both the speed and accuracy of perceptual processing [heitz2014speed].
@@
Recently, a new class of image-computable decision models has sought to fill this gap by explicitly linking sensory encoding with decision dynamics [chen2017seeing, goetschalckx2023computing, rafiei2024neural, jaffe2025image]. These models seek to combine the representational capacity of deep neural networks with temporal accumulation of evidence, providing a functional grounding for the latent decision variable in terms of task-relevant information. Doing so enables models to perform complex tasks while accounting for temporal properties of the decision process. However, existing approaches are often directly fit to human response data, or introduce temporal integration mechanisms without a clear normative justification.
@@ 5 Discussion
We thank Thomas Christie for his thoughtful comments on the manuscript and for the conversations that helped shape and develop this line of work. We also thank Jonathan Pillow, Chris Summerfield, Fabian Sinz, Srini Turaga, Leyla Isik, and Janne Lappalainen for their valuable feedback and discussions during the early conceptualization of this project at the Cajal NeuroAI summer course. Hayden R. Johnson was supported by an FWO grant (G053624N). Anastasia N. Krouglova was supported by an FWO grant (G097022N).

### v2

@@ 1 Introduction
A central task of the brain is to process sensory information in order to guide adaptive behavior, often studied through perceptual decision-making. This process is shaped by several biological constraints. Neurons communicate through discrete action potentials (or spikes), the nervous system is subject to noise at multiple levels [faisal2008noise], and operates under limited metabolic resources [laughlin2001energy, stone2018principles]. Together, these factors impose fundamental limits on both the speed and accuracy of perceptual processing [heitz2014speed].
@@
Recently, a new class of image-computable decision models has sought to fill this gap by explicitly linking sensory encoding with decision dynamics [chen2017seeing, goetschalckx2023computing, cheng2024rtify, rafiei2024neural, jaffe2025image]. These models seek to combine the representational capacity of deep neural networks with temporal accumulation of evidence, providing a functional grounding for the latent decision variable in terms of task-relevant information. Doing so enables models to perform complex tasks while accounting for temporal properties of the decision process. However, many existing approaches are directly fit to human response data, or introduce temporal integration mechanisms without a clear normative justification.
@@ 5 Discussion
We thank Thomas Christie for his thoughtful comments on the manuscript and for the conversations that helped shape and develop this line of work. We also thank Jonathan Pillow, Chris Summerfield, Fabian Sinz, Srini Turaga, Leyla Isik, and Janne Lappalainen for their valuable feedback and discussions during the early development of this project at the Cajal NeuroAI summer course. Hayden R. Johnson was supported by an FWO grant (G053624N). Anastasia N. Krouglova was supported by an FWO grant (G097022N).